# RUBIES: a complete census of the bright and red distant Universe with JWST/NIRSpec

### v1

@@ 3.3 Wavelength and flux calibration
Starting from the robust redshifts measured in the previous section (3.2), we select sources for which the H line and [O iii] doublet fall in the wavelength range of the G395M disperser, . Next, we fit the three emission lines using a custom emission line fitting software that accounts for both the broadening of emission lines by the line spread function (LSF) and the undersampling of the LSF by the NIRSpec detectors. The latter is critical: the NIRSpec LSF for a point source has a width of only pixel (de Graaff et al., 2024a), and fitting, e.g., a Gaussian profile to such an undersampled line could severely over- or underestimate the line flux, because the integrated flux across the pixel differs strongly from the flux density at the mid-point. In order to fit a Gaussian line profile to an undersampled emission line, we first construct a model on an oversampled grid and subsequently integrate the line across the detector pixel.
We assume a single Gaussian line profile for the H and [O iii] doublet, and assume the same kinematics for both line species by fitting for a single velocity dispersion parameter . For a small number of sources with strong outflows this may be inaccurate, but on average we find that these assumptions do not lead to significant residuals. We do not fix the flux ratio of the [O iii] doublet in order to verify that our measurements retrieve the expected theoretical ratio. After constructing the model, we convolve the emission lines with the LSF of an idealised point source (see Appendix A of de Graaff et al., 2024a). This LSF is likely too narrow for many of the spatially-extended sources in the sample, and we therefore do not consider the velocity dispersion measurements themselves to be physically meaningful. We note that we do not use the LSF curves provided in the JWST User Documentation (JDox), as we find these to be too broad for many of our sources and would therefore yield incorrect fluxes.
@@ 4 Science objectives
With approximately 4500 targeted sources spread over a wide area of arcmin2, RUBIES is among the largest spectroscopic programmes performed with JWST/NIRSpec thus far. The combination of low- and medium-resolution spectroscopy allows for a detailed characterisation of the stellar population properties, properties of the interstellar medium, dust and active black holes. The broad range in colour, magnitude and redshift spanned by the RUBIES targets opens up a wealth of opportunities to investigate the growth of galaxies and black holes in the early Universe.
@@ 4.1 Nature of the reddest and brightest high-redshift sources
We detect emission lines in many of these sources, allowing for a precise redshift determination that is difficult to obtain from broad-band photometry alone. We typically find strong H and Paschen line emission indicative of high star formation rates (typically at ; similar to the findings of Barrufet et al., 2024), and a suite of forbidden emission lines in a subset of sources. In conjunction with existing far-infrared and sub-mm constraints from Herschel and ALMA, these measurements place unique constraints on the ISM and dust properties of massive star-forming galaxies at cosmic noon (a first exploration of which is presented in Cooper et al. in prep.). Moreover, by leveraging the full continuum SED, we are able to constrain the stellar population properties and trace the stellar mass growth of the most massive galaxies at (Gottumukkala et al. in prep.).
@@
RUBIES has discovered two of the most extreme such systems to date: an extremely massive () quiescent galaxy at that formed and quenched in the epoch of reionisation (de Graaff et al., 2024b); and a massive () quiescent galaxy at that is a likely progenitor of the massive quiescent galaxies seen at (Weibel et al., 2024a). The analysis of the full population of high-redshift massive quiescent galaxies in RUBIES will set a benchmark for future galaxy formation models and their uncertain models for physical processes such as AGN feedback.
@@
Figure 11: Stellar mass vs. spectroscopic redshift and rest-frame colour, estimated from eazy fitting to PSF-matched photometry of Weibel et al. (2024b) for targets with robust spectroscopic redshifts. Sources that were selected with high priority based on their colour are marked in red. The RUBIES sources span a wide range in redshift, stellar mass and rest-frame colour, with the census sample being predominantly blue and of lower stellar mass. The red sources tend to be massive (), although there is large diversity in the rest-frame colour at the highest stellar masses. We identify a redshift clustering of approximately 15 massive () red galaxies at , which also corresponds to a spatial clustering in the UDS field and coincides with the early massive quiescent galaxy of Glazebrook et al. (2024). The very first Ruby observed, RUBIES-BLAGN-1 at (Wang et al., 2024a), is one of the brightest LRDs discovered thus far, and one of very few LRDs with a MIRI detection in the rest-frame mid-infrared. This single source already demonstrated the complex nature of these systems: broad Balmer () lines are consistent with an actively accreting black hole, but the relatively faint rest-frame mid-infrared emission strongly disfavours the presence of a hot dusty torus that would typically be expected from AGN. RUBIES follow-up of the sources suggested to be in tension with CDM (Labbé et al., 2023) revealed a similarly complex puzzle: broad Balmer lines and narrow [O iii] lines suggest the presence of an AGN; remarkably, however, the SEDs also show Balmer breaks consistent with evolved stellar populations and possibly very high stellar masses at (Wang et al., 2024b).
@@ 4.2 Census of the galaxy population
Star formation histories at . For the continuum-bright sample the PRISM spectra encode critical information on the stellar population properties: the spectra lift the degeneracy between spectral breaks and strong emission lines that affect photometric measurements. In addition, for the brightest sources we even detect stellar absorption features at the PRISM resolution (e.g. de Graaff et al., 2024b). Full spectrum fitting will therefore provide crucial insight into the star formation histories of a diverse population of galaxies at . Moreover, by accounting for the RUBIES selection function, we will also be able to spectroscopically constrain the stellar mass function, which so far has remained restricted to photometric samples at these high redshifts (e.g. Weaver et al., 2023; Harvey et al., 2024; Weibel et al., 2024b; Wang et al., 2024c).
@@
With the diverse population of galaxies probed in RUBIES, we will be able to extend previous studies of the dust attenuation, metallicity and ionisation conditions of the ISM (e.g. Shapley et al., 2023; Sanders et al., 2023; Backhaus et al., 2024) to larger samples, and, critically, to unexplored regions of parameter space. The medium-resolution data additionally reveal the ionised gas kinematics and will allow for a systematic exploration of outflows from star formation and/or AGN (e.g. Xu et al., 2023; Carniani et al., 2024). Furthermore, with the well-quantified selection function of RUBIES we will be able to measure key scaling relations that have so far been limited to relatively small spectroscopic samples or larger photometric samples, such as the star formation main sequence (e.g. Clarke et al., 2024; Rinaldi et al., 2024), or the stellar mass-metallicity relation (e.g. Nakajima et al. 2023; Curti et al. 2024; Lewis et al. in prep.).
@@
We provide a first look into the rest-frame properties of this census sample in Figure 11, which shows targets with robust spectroscopic redshifts. The majority of these sources are at , and span a wide range in stellar mass () and rest-frame colour ( mag). Here, rest-frame properties were estimated by re-running eazy using the PSF-matched photometry from the W24 catalogues, with redshifts fixed to the (robust) spectroscopic redshifts.
@@ 5 Conclusions and data release
RUBIES is a 61-hour Cycle 2 programme with JWST/NIRSpec designed to study the red and bright sources that have been newly discovered with JWST/NIRCam in the EGS and UDS fields. With a total of red objects (), of which are extremely red (), RUBIES provides the first statistical samples of rare objects in the early Universe. Crucially, a carefully constructed census sample of sources is able to place these rare sources in the context of the broader galaxy population at : this census sample holds immense legacy value, and is enabled by a custom-developed mask design strategy for NIRSpec that samples the full parent population and relies solely on the measured F444W magnitudes, colours, and photometric redshifts of sources.
Obtaining such a large spectroscopic sample of red sources is crucial, as this population is highly heterogeneous. We find that the red sources span a wide redshift range, from , and show diverse spectral properties that do not correlate trivially with redshift, magnitude or colour alone. In comparison to the full galaxy population, the red sources are among the most massive systems at all redshifts and therefore contribute significantly to the stellar mass and star formation rate density in the early Universe.
@@ Acknowledgements.
We thank the CEERS and PRIMER teams for making their imaging data publicly available immediately. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1345, #1837 #2234, #2279, #2514, #2750, #3990 and #4233. Support for program #4233 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #562. The Cosmic Dawn Center is funded by the Danish National Research Foundation (DNRF) under grant #140. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. Support for this work for RPN was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.References
@@
- Barrufet et al. (2024) Barrufet, L., Oesch, P., Marques-Chaves, R., et al. 2024, arXiv e-prints, arXiv:2404.08052
@@
- Bunker et al. (2023) Bunker, A. J., Cameron, A. J., Curtis-Lake, E., et al. 2023, arXiv e-prints, arXiv:2306.02467
@@
- Cameron et al. (2024) Cameron, A. J., Katz, H., Witten, C., et al. 2024, MNRAS[arXiv:2311.02051]
@@
- Chisholm et al. (2024) Chisholm, J., Berg, D. A., Endsley, R., et al. 2024, arXiv e-prints, arXiv:2402.18643
- Clarke et al. (2024) Clarke, L., Shapley, A. E., Sanders, R. L., et al. 2024, arXiv e-prints, arXiv:2406.05178
@@
- de Graaff et al. (2024a) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024a, A&A, 684, A87
- de Graaff et al. (2024b) de Graaff, A., Setton, D. J., Brammer, G., et al. 2024b, arXiv e-prints, arXiv:2404.05683
- D’Eugenio et al. (2024) D’Eugenio, F., Cameron, A. J., Scholtz, J., et al. 2024, arXiv e-prints, arXiv:2404.06531
@@
- Donnan et al. (2024) Donnan, C. T., McLure, R. J., Dunlop, J. S., et al. 2024, arXiv e-prints, arXiv:2403.03171
@@
- Fujimoto et al. (2023a) Fujimoto, S., Arrabal Haro, P., Dickinson, M., et al. 2023a, ApJ, 949, L25
- Fujimoto et al. (2023b) Fujimoto, S., Wang, B., Weaver, J., et al. 2023b, arXiv e-prints, arXiv:2308.11609
@@
- Harvey et al. (2024) Harvey, T., Conselice, C., Adams, N. J., et al. 2024, arXiv e-prints, arXiv:2403.03908
- Heintz et al. (2024) Heintz, K. E., Brammer, G. B., Watson, D., et al. 2024, arXiv e-prints, arXiv:2404.02211
- Herard-Demanche et al. (2023) Herard-Demanche, T., Bouwens, R. J., Oesch, P. A., et al. 2023, arXiv e-prints, arXiv:2309.04525
@@
- Labbe et al. (2023) Labbe, I., Greene, J. E., Bezanson, R., et al. 2023, arXiv e-prints, arXiv:2306.07320
@@
- Maseda et al. (2024) Maseda, M. V., de Graaff, A., Franx, M., et al. 2024, arXiv e-prints, arXiv:2403.05506
@@
- Price et al. (2023) Price, S. H., Suess, K. A., Williams, C. C., et al. 2023, arXiv e-prints, arXiv:2310.02500
@@
- Rinaldi et al. (2024) Rinaldi, P., Navarro-Carrera, R., Caputi, K. I., et al. 2024, arXiv e-prints, arXiv:2406.13554
@@
- Weibel et al. (2024a) Weibel, A., de Graaff, A., Setton, D. J., Miller, T. B., & Oesch, P. A. 2024a, arXiv e-prints, arXiv:2409.03829
@@
- Xiao et al. (2023) Xiao, M., Oesch, P., Elbaz, D., et al. 2023, arXiv e-prints, arXiv:2309.02492

### v2

@@ 3.3 Wavelength and flux calibration
Starting from the robust redshifts measured in the previous section (3.2), we select sources for which the H line and [O iii] doublet fall in the wavelength range of the G395M disperser, . Next, we fit the three emission lines using a custom emission line fitting software that accounts for both the broadening of emission lines by the line spread function (LSF) and the undersampling of the LSF by the NIRSpec detectors. The latter is critical, as the NIRSpec LSF for a point source has a width of only pixel (de Graaff et al., 2024), and fitting, e.g., a Gaussian profile to such an undersampled line could severely over- or underestimate the line flux: because the flux density profile is highly non-linear across the pixel, the integrated flux of the profile across the pixel differs from the flux obtained by simply evaluating the profile at the mid-point of the pixel. In order to robustly fit a Gaussian line profile to the undersampled NIRSpec data, we therefore first construct the Gaussian emission line model on a fine wavelength grid (a factor 5 higher than the NIRSpec wavelength sampling) and subsequently integrate the model with a Riemann sum.
We assume a single Gaussian line profile for the H and [O iii] doublet, and assume the same kinematics for both line species by fitting for a single velocity dispersion parameter . For a small number of sources with strong outflows this may be inaccurate, but on average we find that these assumptions do not lead to significant residuals. We do not fix the flux ratio of the [O iii] doublet in order to verify that our measurements retrieve the expected theoretical ratio. After constructing the model, we convolve the emission lines with the LSF of an idealised point source (see Appendix A of de Graaff et al., 2024). This LSF is likely too narrow for many of the spatially-extended sources in the sample, and we therefore do not consider the velocity dispersion measurements themselves to be physically meaningful. We note that we do not use the LSF curves provided in the JWST User Documentation (JDox), as we find these to be too broad for many of our sources and would therefore yield incorrect fluxes.
@@ 4 Science objectives
With 4444 targeted sources spread over a wide area of arcmin2, RUBIES is among the largest spectroscopic programmes performed with JWST/NIRSpec thus far. The combination of low- and medium-resolution spectroscopy allows for a detailed characterisation of the stellar population properties, properties of the interstellar medium, dust and active black holes. The broad range in colour, magnitude and redshift spanned by the RUBIES targets opens up a wealth of opportunities to investigate the growth of galaxies and black holes in the early Universe.
@@ 4.1 Nature of the reddest and brightest high-redshift sources
We detect emission lines in many of these sources, allowing for a precise redshift determination that is difficult to obtain from broad-band photometry alone. We typically find strong H and Paschen line emission indicative of high star formation rates (typically at ; similar to the findings of Barrufet et al., 2025), and a suite of forbidden emission lines in a subset of sources. In conjunction with existing far-infrared and sub-mm constraints from Herschel and ALMA, these measurements place unique constraints on the ISM and dust properties of massive star-forming galaxies at cosmic noon (a first exploration of which is presented in Cooper et al. 2024). Moreover, by leveraging the full continuum SED, we are able to constrain the stellar population properties and trace the stellar mass growth of the most massive galaxies at (Gottumukkala et al. in prep.).
@@
RUBIES has discovered two of the most extreme such systems to date: an extremely massive () quiescent galaxy at that formed and quenched in the epoch of reionisation (de Graaff et al., 2025); and a massive () quiescent galaxy at that is a likely progenitor of the massive quiescent galaxies seen at (Weibel et al., 2024a). The analysis of the full population of high-redshift massive quiescent galaxies in RUBIES will set a benchmark for future galaxy formation models and their uncertain models for physical processes such as AGN feedback.
@@
Figure 11: Stellar mass vs. spectroscopic redshift and rest-frame colour, estimated from eazy fitting to PSF-matched photometry of Weibel et al. (2024b) for targets with robust spectroscopic redshifts. Sources that were selected with high priority based on their colour are marked in red. The RUBIES sources span a wide range in redshift, stellar mass and rest-frame colour, with the census sample being predominantly blue and of lower stellar mass. The red sources tend to be massive (), although there is large diversity in the rest-frame colour at the highest stellar masses. We identify a redshift clustering of approximately 15 massive () red galaxies at , which also corresponds to a spatial clustering in the UDS field and coincides with the early massive quiescent galaxy of Glazebrook et al. (2024). The very first ‘Ruby’ (Priority 1 target) observed, RUBIES-BLAGN-1 at (Wang et al., 2024a), is one of the brightest LRDs discovered thus far, and one of very few LRDs with a MIRI detection in the rest-frame mid-infrared. This single source already demonstrated the complex nature of these systems: broad Balmer () lines are consistent with an actively accreting black hole, but the relatively faint rest-frame mid-infrared emission strongly disfavours the presence of a hot dusty torus that would typically be expected from AGN. RUBIES follow-up of the sources suggested to be in tension with CDM (Labbé et al., 2023) revealed a similarly complex puzzle: broad Balmer lines and narrow [O iii] lines suggest the presence of an AGN; remarkably, however, the SEDs also show Balmer breaks consistent with evolved stellar populations and possibly very high stellar masses at (Wang et al., 2024b).
@@ 4.2 Census of the galaxy population
Star formation histories at . For the continuum-bright sample the PRISM spectra encode critical information on the stellar population properties: the spectra lift the degeneracy between spectral breaks and strong emission lines that affect photometric measurements. In addition, for the brightest sources we even detect stellar absorption features at the PRISM resolution (e.g. de Graaff et al., 2025). Full spectrum fitting will therefore provide crucial insight into the star formation histories of a diverse population of galaxies at . Moreover, by accounting for the RUBIES selection function, we will also be able to spectroscopically constrain the stellar mass function, which so far has remained restricted to photometric samples at these high redshifts (e.g. Weaver et al., 2023; Harvey et al., 2025; Weibel et al., 2024b; Wang et al., 2024c).
@@
With the diverse population of galaxies probed in RUBIES, we will be able to extend previous studies of the dust attenuation, metallicity and ionisation conditions of the ISM (e.g. Shapley et al., 2023; Sanders et al., 2023; Backhaus et al., 2024) to larger samples, and, critically, to unexplored regions of parameter space. The medium-resolution data additionally reveal the ionised gas kinematics and will allow for a systematic exploration of outflows from star formation and/or AGN (e.g. Xu et al., 2023; Carniani et al., 2024). Furthermore, with the well-quantified selection function of RUBIES we will be able to measure key scaling relations that have so far been limited to relatively small spectroscopic samples or larger photometric samples, such as the star formation main sequence (e.g. Clarke et al., 2024; Rinaldi et al., 2025), or the stellar mass-metallicity relation (e.g. Nakajima et al. 2023; Curti et al. 2024; Lewis et al. in prep.).
@@
We provide a first look into the rest-frame properties of this census sample in Figure 11, which shows targets with robust spectroscopic redshifts. The majority of these sources are at , and span a wide range in stellar mass () and rest-frame colour ( mag). Here, rest-frame properties were estimated by cross-matching our targets with the PSF-matched photometry from the W24 catalogues and re-running eazy with redshifts fixed to the (robust) spectroscopic redshifts.
@@ 5 Conclusions and data release
RUBIES is a 60-hour Cycle 2 programme with JWST/NIRSpec designed to study the red and bright sources that have been newly discovered with JWST/NIRCam in the EGS and UDS fields. With a total of red objects (), of which are extremely red (), RUBIES provides the first statistical samples of rare objects in the early Universe. Crucially, a carefully constructed census sample of sources is able to place these rare sources in the context of the broader galaxy population at : this census sample holds immense legacy value, and is enabled by a custom-developed mask design strategy for NIRSpec that samples the full parent population and relies solely on the measured F444W magnitudes, colours, and photometric redshifts of sources.
Obtaining such a large spectroscopic sample of red sources is crucial, as this population is highly heterogeneous. We find that the red sources span a wide redshift range, from , and show diverse spectral properties that do not correlate trivially with redshift, magnitude or colour alone. In comparison to the full galaxy population, the red sources are among the most massive systems at all redshifts and therefore could possibly contribute significantly to the stellar mass and star formation rate density in the early Universe.
@@ Acknowledgements.
We thank the CEERS and PRIMER teams for making their imaging data publicly available immediately. This work is based on observations made with the NASA/ESA/CSA James Webb Space Telescope. The data were obtained from the Mikulski Archive for Space Telescopes at the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127 for JWST. These observations are associated with programs #1345, #1837 #2234, #2279, #2514, #2750, #3990 and #4233. Support for program #4233 was provided by NASA through a grant from the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS 5-03127. REH acknowledges support by the German Aerospace Center (DLR) and the Federal Ministry for Economic Affairs and Energy (BMWi) through program 50OR2403 ‘RUBIES’. This research was supported by the International Space Science Institute (ISSI) in Bern, through ISSI International Team project #562. The Cosmic Dawn Center is funded by the Danish National Research Foundation (DNRF) under grant #140. This work has received funding from the Swiss State Secretariat for Education, Research and Innovation (SERI) under contract number MB22.00072, as well as from the Swiss National Science Foundation (SNSF) through project grant 200020_207349. Support for this work for RPN was provided by NASA through the NASA Hubble Fellowship grant HST-HF2-51515.001-A awarded by the Space Telescope Science Institute, which is operated by the Association of Universities for Research in Astronomy, Incorporated, under NASA contract NAS5-26555.References
@@
- Arrabal Haro et al. (2023a) Arrabal Haro, P., Dickinson, M., Finkelstein, S. L., et al. 2023a, ApJ, 951, L22
- Arrabal Haro et al. (2023b) Arrabal Haro, P., Dickinson, M., Finkelstein, S. L., et al. 2023b, Nature, 622, 707
@@
- Barrufet et al. (2025) Barrufet, L., Oesch, P. A., Marques-Chaves, R., et al. 2025, MNRAS, 537, 3453
@@
- Bunker et al. (2024) Bunker, A. J., Cameron, A. J., Curtis-Lake, E., et al. 2024, A&A, 690, A288
@@
- Cameron et al. (2024) Cameron, A. J., Katz, H., Witten, C., et al. 2024, MNRAS, 534, 523
@@
- Chisholm et al. (2024) Chisholm, J., Berg, D. A., Endsley, R., et al. 2024, MNRAS, 534, 2633
- Clarke et al. (2024) Clarke, L., Shapley, A. E., Sanders, R. L., et al. 2024, ApJ, 977, 133
- Cooper et al. (2024) Cooper, O. R., Brammer, G., Heintz, K. E., et al. 2024, arXiv e-prints, arXiv:2410.08387
@@
- de Graaff et al. (2024) de Graaff, A., Rix, H.-W., Carniani, S., et al. 2024, A&A, 684, A87
- de Graaff et al. (2025) de Graaff, A., Setton, D. J., Brammer, G., et al. 2025, Nature Astronomy, 9, 280
- D’Eugenio et al. (2025) D’Eugenio, F., Cameron, A. J., Scholtz, J., et al. 2025, ApJS, 277, 4
@@
- Donnan et al. (2024) Donnan, C. T., McLure, R. J., Dunlop, J. S., et al. 2024, MNRAS, 533, 3222
@@
- Finkelstein et al. (2025) Finkelstein, S. L., Bagley, M. B., Arrabal Haro, P., et al. 2025, arXiv e-prints, arXiv:2501.04085
@@
- Fujimoto et al. (2023) Fujimoto, S., Arrabal Haro, P., Dickinson, M., et al. 2023, ApJ, 949, L25
- Fujimoto et al. (2024) Fujimoto, S., Wang, B., Weaver, J. R., et al. 2024, ApJ, 977, 250
@@
- Harvey et al. (2025) Harvey, T., Conselice, C. J., Adams, N. J., et al. 2025, ApJ, 978, 89
- Heintz et al. (2025) Heintz, K. E., Brammer, G. B., Watson, D., et al. 2025, A&A, 693, A60
- Herard-Demanche et al. (2025) Herard-Demanche, T., Bouwens, R. J., Oesch, P. A., et al. 2025, MNRAS, 537, 788
@@
- Labbe et al. (2025) Labbe, I., Greene, J. E., Bezanson, R., et al. 2025, ApJ, 978, 92
@@
- Maseda et al. (2024) Maseda, M. V., de Graaff, A., Franx, M., et al. 2024, A&A, 689, A73
@@
- Price et al. (2025) Price, S. H., Suess, K. A., Williams, C. C., et al. 2025, ApJ, 980, 11
@@
- Rinaldi et al. (2025) Rinaldi, P., Navarro-Carrera, R., Caputi, K. I., et al. 2025, ApJ, 981, 161
@@
- Weibel et al. (2024a) Weibel, A., de Graaff, A., Setton, D. J., et al. 2024a, arXiv e-prints, arXiv:2409.03829
@@
- Xiao et al. (2024) Xiao, M., Oesch, P. A., Elbaz, D., et al. 2024, Nature, 635, 311